A local Dunkin Donuts shop reported that it's sales have increased exactly 22% per year for the last 2 years. This year's sales were $82,500. What were Dunkin Donuts sales 2 years ago? Round each years' sales to the nearest dollar. Answer by lwsshak3(11628) 
 A local Dunkin Donuts shop reported that it's sales have increased exactly 22% per year for the last 2 years. This year's sales were $82,500. What were Dunkin Donuts sales 2 years ago? Round each years' sales to the nearest dollar.
**
let x=1st year sales
x+.22x=1.22x=2nd year sales
(1.22x)+(1.22x)*.22=82500
(1.22x)+.2684x=82500
1.4884x=82500
x=82500/1.4884
x=55429
1.22x=67623
..
1st year sales=$55,429
2nd year sales=$67,623